OpenStudy (anonymous):

The length of a rectangular garden is 3 yard more than twice its width. the perimeter of the garden is 36yd. what are the width and length of the garden?

OpenStudy (anonymous):

If the width is w, what's the length (in terms of w)?

OpenStudy (anonymous):

x + x + 2x + 3 + 2x +3 = 36 the 2x's are for the "twice the width" part and the two 3's are for the 3yds more part. then combine like terms 6x+6 = 36 minus the 6 over 6x = 36-6 6x = 30 then just finish out with algebra x = 30/6 x = 5yds

OpenStudy (anonymous):

since the length is 3 yards more than twice the width then l=3+2x and x is the width. You then add 2(3+2x)+2(x) since there are two widths and length and set that equal to 36. 6+4x+2x=36. 6x+6=36 6x=30 x=5 plug in your value into the length and width equations (3+2x) and (x) to find length and width